how to make calculation very strong in pcm.......

Grade:12

2 Answers

AKASH GOYAL AskiitiansExpert-IITD
13 years ago

Dear Syed

learn the squares vales at least upto 30, square roots upto 20 and learn the fraction values like 1/3=0.33, 1/6=0.166, 1/7=0.143 etc upto 1/20 at least

it will make your calculations really fast without calculators

 

 All the best.                                                          

AKASH GOYAL

AskiitiansExpert-IIT Delhi

 

Please feel free to post as many doubts on our discussion forum as you can. We are all IITians and here to help you in your IIT JEE preparation.

Win exciting gifts by answering the questions on Discussion Forum. So help discuss any query on askiitians forum and become an Elite Expert League askiitian.

Now you score 5+15 POINTS by uploading your Pic and Downloading the Askiitians Toolbar  respectively : Click here to download the toolbar..

 

 

Anupam Dalmia
13 years ago

Practise makes a man perfect.

 

PRACTISE a lot

 

u must remember all the formulae and constants with their values

u can put up a chart in your room to remember those.